OpenStudy (anonymous):

Use the function to answer the question. f(x) = 4|x + 1| - 6 What is the vertex of the graph of f(x)?

OpenStudy (anonymous):

x is greater than or equal to zero what i did was plugged in a few numbers (2,1,0,-1,-2) and then checked the answer. if you practice doing that for a while it's easy to do

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

also u should look it up before u post it on Os,Mostly Yahoo answer can help u

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

hoped this helped

OpenStudy (anonymous):

It did, but I can't figure it out still! Lol

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

Given f(x) = 6 - x (I assume this since [x] means something which I do not think you want it to mean) f(-1.3) = 6 - 1.3 = 4.7, which isn't one of your options. So I can only deduce you REALLY meant f(x) = 6( -|x|) Thus f(-1.3) = 6(-1.3) = -7.8 In the future please state your functions more clearly. ________ Pay this no attention, original poster_________ And for anyone interested, as the question was worded. Given f(x) = 6 - [x], f(-1.3) = 5.

OpenStudy (anonymous):

lol live

OpenStudy (anonymous):

U will need to Talk an expert let me connect *elevator music*

OpenStudy (anonymous):

live gave u the wrong answer

OpenStudy (jdoe0001):

what value of "x" makes \(\bf |x + 1|\ = 0\quad ?\)

OpenStudy (anonymous):

ur wrong

OpenStudy (anonymous):

what value of "x" makes Ix+1I= 0

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

1x...?

OpenStudy (anonymous):

oh that 1 is easy alex. 1x1=1

OpenStudy (anonymous):

understood?

OpenStudy (jdoe0001):

AlexLovesGod so.... what value of "x" makes the absolute expression equals to 0? what do you think?

OpenStudy (anonymous):

AlexLovesGod so.... what value of "x" makes the absolute expression equals to 0? what do you think?

OpenStudy (anonymous):

I AM SO CONFUSED!!!! why is live copying jdoe and I am lost

OpenStudy (anonymous):

please no questions

OpenStudy (jhannybean):

@live10000000 let @jdoe0001 help OP.

OpenStudy (jdoe0001):

AlexLovesGod so... what do you think? |x + 1 | = 0 <--- what value of "x" will give us that?

OpenStudy (anonymous):

Please you are not involved in this question

OpenStudy (anonymous):

my face hurts

OpenStudy (jdoe0001):

face hurts? that's a first one I've heard hehe usually is my head hurts... or my cheeks...